An object is moving around the unit circle with paI just need help with question F
Answer #1

\underline{Solution}:

From problem (a)

Area\, \, of\, \, a\, \, triangle\,\, a(t)=\frac{1}{sin2t}

Given\,\, a(t)=2

\Rightarrow \frac{1}{sin2t}=2

\Rightarrow sin2t=\frac{1}{2}

\Rightarrow sin2t=\frac{1}{2}=sin\frac{\pi }{6}................................(1)

\Rightarrow Princple\, value\, (\alpha )=\frac{\pi }{6}

General\, solution\, of\, (1)\, is\,\, 2t=n\pi +(-1)^{n}\alpha

\Rightarrow 2t=n\pi +(-1)^{n}\frac{\pi }{6}

\Rightarrow t=\frac{n\pi}{2} +(-1)^{n}\frac{\pi }{12};\, \, \, n\epsilon Z

If\, n=0\Rightarrow t=0 +\frac{\pi }{12}\Rightarrow t=\frac{\pi }{12}=15^{0}

If\, n=1\Rightarrow t=\frac{\pi}{2}-\frac{\pi }{12}\Rightarrow t=\frac{5\pi }{12}=75^{0}

If\, n=2\Rightarrow t=\pi+\frac{\pi }{12}\Rightarrow t=\frac{13\pi }{12}=195^{0}

If\, n=-1\Rightarrow t=-\frac{\pi }{2}-\frac{\pi }{12}\Rightarrow t=-\frac{7\pi }{12}=-105^{0}

Since\, \, 0< t< \frac{\pi }{2}\, \, \left [ i.e.,0< t< 90^{0} \right ]

\Rightarrow t=\frac{\pi}{12},\, \, t=\frac{5\pi }{12}

\therefore Smalest \, \, t=\frac{\pi}{12}\, \, and\, \, Largest\, \, t=\frac{5\pi }{12}

.........................................................................................................................................